# Attractors for Singular-Degenerate Porous Medium Type Equations Arising in Models for Biofilm Growth

### v1

@@ Abstract
We investigate the long-time behaviour of solutions of a class of singular-degenerate porous medium type equations in bounded Lipschitz domains with mixed Dirichlet-Neumann boundary conditions. The existence of global attractors is shown under very general assumptions. Assuming, in addition, that solutions are globally Hölder continuous and the reaction terms satisfy a suitable sign condition in the vicinity of the degeneracy, we also prove the existence of an exponential attractor, which, in turn, yields the finite fractal dimension of the global attractor. Moreover, we extend the results for scalar equations to systems where the degenerate equation is coupled to a semilinear reaction-diffusion equation. The study of such systems is motivated by models for biofilm growth.
@@ 1.1 Singular-degenerate porous medium type equations
complemented with a suitable initial data and mixed Dirichlet-Neumann boundary conditions, where $T>0$ and $\Omega\subset\mathbb{R}^{n},n\geq 1$, is a bounded Lipschitz domain. The solution $u$ assumes its values in the interval $(-1,1)$, $f:\Omega\times(-1,1)\to\mathbb{R}$ is continuous and $\phi:(-1,1)\to\mathbb{R}$ is strictly increasing with a degeneracy $\phi^{\prime}(0)=0$ and singularities $\phi(\pm 1)=\pm\infty$. Since the degeneracy in $u=0$ is of the same type as in the porous medium equation, $u_{t}=\Delta u^{m},m>1$, we call (1.1) a singular-degenerate equation of porous medium type. Porous medium type equations are used to model various natural phenomena such as the flow of an ideal gas through a porous medium ([28]), nonlinear heat transfer ([34]) and the spreading of certain biological populations ([18]).
@@
equipped with suitable initial values and mixed Dirichlet-Neumann boundary data. Such systems arise in models for the growth of bacterial biofilms. Biofilms are dense aggregations of microorganisms that form in moist environments. In a biofilm, cells are encased in a slimy matrix of extracellular polymeric substances and adhere to one another and usually to surfaces. A reaction-diffusion model that predicts heterogeneous spatial structures of fully grown biofilms on the mesoscale was introduced and numerically studied by Eberl et al. in [7]. It is a system of a quasilinear reaction-diffusion equation for the biomass density $M$ and a semilinear reaction-diffusion equation for the nutrient concentration $C$. The equations are coupled via Monod reaction functions modeling biomass production and nutrient consumption. For the dimensionless variables $C$ and $M$, which are normalized, respectively, with respect to the bulk concentration and the maximum biomass density, the model reads as follows:
@@
[…]
we observe that the biofilm-growth model (1.3) is a particular case of the system (1.2) that we consider. Moreover, in simulation studies, $\Omega$ is typically rectangular and mixed Dirichlet-Neumann boundary conditions are imposed. More specifically, homogeneous Neumann boundary conditions for $C$ and $M$ are assumed on the bottom part of the boundary, due to the impermeability of the surface to nutrients and biomass, and also on the lateral boundaries. Nutrients are added through the top boundary, which is described by Dirichlet boundary conditions for $C$, while homogeneous Dirichlet boundary conditions are imposed for $M$. Initially, there are small pockets of biomass of density $M_{0}<1$ on the bottom part of the boundary, resembling the initial biofilm colonies on the surface, while $M_{0}$ vanishes everywhere else in the domain. The nutrient concentration $C_{0}$ is set to the bulk concentration everywhere. Therefore, the results of this paper apply to the biofilm growth model (1.3) with such initial data and mixed Dirichlet-Neumann boundary conditions on bounded Lipschitz domains.
@@ 1.2 Global and exponential attractors: previous results
The first existence proof for exponential attractors in [9] was non-constructive, based on the squeezing property of the semigroup and restricted to a Hilbert space setting. Efendiev et al. [10] provided an alternative, explicit construction of exponential attractors for semigroups in Banach spaces based on the decomposition the semigroup into a compact and contracting part. In [5], Czaja and Efendiev extended this construction and the existence proof for exponential attractors to a setting that can be applied to degenerate parabolic equations. It covers a noteworthy smoothing estimate formulated in suitable function spaces related to pieces of trajectories of the semigroup and not in the phase space. Efendiev and Zelik further generalized this approach allowing for estimates and function spaces that depend on points in a compact, positively invariant set to construct exponential attractors for equations of porous medium type, see [11, Remark 3.2].
@@
in $(0,\infty)\times\mathbb{R}^{n}$ and proved, by applying the splitting method, that the generated semigroup possesses a global attractor in the weighted space $L^{1}(\mathbb{R}^{n};\rho(x)dx)$ where the weight $\rho\in L^{1}(\mathbb{R}^{n})$ goes to zero at infinity. To prove the asymptotic compactness of the semigroup for our problem we use arguments from this paper. Andreu et al. [1] proved the existence of a global attractor in $L^{\infty}(\Omega)$ for (1.4) with $g=0$ in a bounded smooth domain $\Omega\subset\mathbb{R}^{n}$ with nonlinear Neumann boundary conditions under stronger assumptions on the nonlinearities $\phi$ and $f$. Moreover, Carvalho et al. [3] considered (1.4) with $\phi(u)=|u|^{m}$, $m>1$, and $g=0$ in bounded smooth domains with homogeneous Dirichlet boundary conditions and a globally Lipschitz continuous function $f$ and proved the existence of a global attractor in $H^{-1}(\Omega)$ based on the theory of maximal monotone operators. Finally, in the aforementioned paper [12], the authors showed that the biofilm growth model (1.3) in a smooth bounded domain $\Omega$ generates a Lipschitz continuous semigroup in $L^{1}(\Omega)$ that possesses a global attractor.
@@ 1.3 Aim and outline
Our aim is to prove the existence of global and exponential attractors for problems (1.1) and (1.2) under suitable structural assumptions. The existence of exponential attractors ensures the existence and finite fractal dimension of the global attractor. However, we state a separate existence theorem for global attractors as it holds under much weaker assumptions. It generalizes the results in [12] on the existence of a global attractor for the biofilm model (1.3). To construct exponential attractors we additionally need a sign condition on $f$ at zero and the global Hölder continuity of solutions, similarly as in [11]. We simplify the latter proof, where first the finite fractal dimension of the global attractor was shown and then the exponential attractor was constructed, by using an abstract existence theorem for exponential attractors, see [11, Remark 3.2] and [5]. The finite fractal dimension of the global attractor is then a consequence of the existence of an exponential attractor. Our results are formulated for a broad class of degenerate problems under mixed Dirichlet-Neumann boundary data in bounded Lipschitz domains. To the best of our knowledge, they are novel in the context of global and exponential attractors. We remark that system (1.2) is not restricted to biofilm models. Such systems can also be used to describe other phenomena with sharp interfaces propagating at a finite speed. Moreover, our results easily extend to, e.g., the model for quorum sensing induced detachment in biofilms [15] which was analytically studied in [16]. It is an extension of the biofilm growth model (1.3) involving one singular-degenerate equation and three semi-linear reaction diffusion equations for dissolved substrates.
We organize the paper as follows: In Section 2, we introduce the functional setting and state the main results on the existence of global and exponential attractors. In Section 3, we first prove the existence of a global attractor for (1.1) in a suitable phase space endowed with the $L^{1}(\Omega)$-topology. To this end, we show the existence of a bounded absorbing set and the asymptotic compactness of the trajectories, and then extend the result to the coupled system (1.2). In Section 4, under the key assumptions of global Hölder continuity of solutions and appropriate sign conditions on the partial derivative of the nonlinearity $f$ in the vicinity of the degeneracy, we prove the existence of exponential attractors for (1.1). As a consequence, the global attractor has finite fractal dimension. We also extend the existence result for exponential attractors for the coupled system (1.2).
@@ 2.1 Well-Posedness Results
We recall the well-posedness results in [21] and [22] for the scalar equation (1.1) and the coupled system (1.2), and introduce the functional setting for the associated semigroups. First, we consider the following initial-boundary value problem for (1.1) with mixed Dirichlet-Neumann boundary conditions,
$\left\{\begin{array}[]{l}u_{t}=\Delta\phi(u)+f(\cdot,u)\ \ \text{in}\ (0,T]\times\Omega,\\ u=u_{0}\ \ \text{in}\ \{0\}\times\Omega,\\ \phi(u)=\phi(u^{D})\ \ \text{on}\ (0,T]\times\Gamma,\\ \partial_{\nu}\phi(u)=0\ \ \text{on}\ (0,T]\times\partial\Omega/\Gamma,\end{array}\right.$ (2.1)
where $T>0$, $\Omega\subset\mathbb{R}^{n},n\geq 1,$ is a bounded Lipschitz domain with boundary $\partial\Omega$, and $\Gamma\subset\partial\Omega$ is measurable with Hausdorff measure $H^{n-1}(\Gamma)>0$. Let $\Omega_{T}=(0,T]\times\Omega$ and $I=(-1,1)$.
@@
$\phi\ \text{is surjective,}$, (2.3)
and
$\left\{\begin{array}[]{ll}\phi\ \text{is piece-wise continuously differentiable,}\\ \phi^{\prime}(0)=0\ \text{and}\ \phi\ \text{is\ convex on}\ [0,1)\ \text{and}\ \text{concave on}\ (-1,0].\end{array}\right.$ (2.6)
@@
We assume that the Dirichlet data satisfies
$u^{D}:\Omega\to I\ \text{is measurable and}\ \phi(u^{D})\in L^{\infty}(\Omega)\cap H^{1}(\Omega),$ (2.8)
and the initial data
@@
[…]
Now, we introduce the functional setting for problem (2.1). By (2.2) and (2.8), we have spatial regularity of $u^{D}$, and therefore, we can formulate (2.1)3italic-(2.1subscriptitalic-)$_{3}$ in trace sense. We consider the closed subspace
$V=\{w\in H^{1}(\Omega):Tr\ w=0\ a.e.\ on\ \Gamma\}$
of $H^{1}(\Omega)$ and denote its dual by $V^{*}$. The solutions need to satisfy $\phi(u)\in\phi(u^{D})+L^{2}(0,T;V)$ to ensure that $\phi(u)=\phi(u^{D})$ a.e. on $\Gamma$.
@@
$|z|\leq\delta\Phi(z)+\beta(\frac{1}{\delta}),\quad z\in I,\delta>0,$ (2.10)
@@
Definition 2.1.
A measurable function $u:\Omega_{T}\to I$ is a weak solution of (2.1) if $u\in H^{1}(0,T;V^{*})$ with weak derivative $u_{t}$, $u(0)=u_{0}\in V^{*}$ and $\phi(u)\in\phi(u^{D})+L^{2}(0,T;V)$ that satisfies the scalar equation (2.1)1italic-(2.1subscriptitalic-)$_{1}$ in distributional sense, i.e. the identity
[…]
@@
By Definition 2.1, it follows that $\Phi(u)\in L^{\infty}(0,T;L^{1}(\Omega))$ (see [21], Corollary 3.15. Then, (2.7) and the essential boundedness of $f(\cdot,0)$ imply that $f(\cdot,u)\in L^{\infty}(0,T;L^{2}(\Omega))$ so that the right-hand side of (2.11) is well-defined.
We formulate the following well-posedness result for the scalar problem (2.1), which follows from [[21], Theorems 2.6–2.7].
@@
$\lVert\Phi(u;u^{D})\rVert_{L^{\infty}(0,T;L^{1}(\Omega))}+\lVert\nabla\phi(u)\rVert_{L^{2}(\Omega_{T})}^{2}$
$\leq C\left(\lVert\Phi(u_{0};u^{D})\rVert_{L^{1}(\Omega)}+\lVert\nabla\phi(u^{D})\rVert_{L^{2}(\Omega_{T})}+\lVert f(\cdot,u)\rVert_{L^{2}(\Omega_{T})}\lVert\phi(u)-\phi(u^{D})\rVert_{L^{2}(\Omega_{T})}\right),$ (2.12)
where $\Phi(z;\bar{z}):=\int\limits_{\bar{z}}^{z}\phi(\tilde{z})-\phi(\bar{z})d\tilde{z}$ and $C\geq 0$ some constant.
The solution depends continuously on initial data. More specifically, if $u$ and $\tilde{u}$ are two solutions of (2.1) corresponding to the functions $f$, $\tilde{f}$ satisfying (2.7), the same Dirichlet data $u^{D}$ satisfying (2.8) and initial data $u_{0}$, $\tilde{u_{0}}$ satisfying (2.9), then the following $L^{1}$-contraction holds,
[…]
@@
Moreover, if $f$ is also bounded and $\phi(u_{0})\in L^{\infty}(\Omega)$, then $\phi(u)$ is bounded by a constant depending on $\Omega$ and the bounds of $f$ and $\phi(u_{0})$. If the condition (2.6) holds, then we have $u\in C([0,T];L^{1}(\Omega))$. If both assumptions hold and $\phi(u_{0})\in\phi(u^{D})+V$, then $\phi(u)\in H^{1}(0,T;L^{2}(\Omega))$.
Remark 2.5.
In Theorem 2.1, $\int\limits_{\Omega}\Phi(u;u^{D})$ denotes the relative energy functional of $u$ relative to $u^{D}$. The estimate (2.12) can be formulated in terms of the absolute energy, instead of the relative energy, as follows:
$\lVert\Phi(u)\rVert_{L^{\infty}(0,T;L^{1}(\Omega))}+\lVert\nabla\phi(u)\rVert_{L^{2}(\Omega_{T})}^{2}\leq C,$ (2.14)
for some constant $C\geq 0$ (cf. [21], Remark 2.8). Indeed, by using
$\Phi(u;u^{D})=\Phi(u)-\Phi(u^{D})-\phi(u^{D})(u-u^{D})$
and the inequality $|\phi(u^{D})||u|\leq\frac{1}{2}\Phi(u)+C$ which is obtained by (2.10) with $\delta=\frac{1}{2}\lVert\phi(u^{D})\rVert_{L^{\infty}(\Omega)}^{-1}$, we find
$\Phi(u)\leq C\left(\Phi(u;u^{D})+1\right).$ (2.15)
On the other hand, by Young’s and Poincare’s inequality, the term on the right-hand side of (2.12), including the reaction function, can be estimated by
$C\left(\lVert f(\cdot,u)\rVert_{L^{2}(\Omega_{T})}^{2}+1\right)+\frac{1}{2}\lVert\nabla\phi(u)\rVert_{L^{2}(\Omega_{T})}^{2}.$
Then, since the continuity of $f$ with respect to the last argument and the boundedness of $u$ imply that
$\lVert f(\cdot,u)\rVert_{L^{2}(\Omega_{T})}^{2}\leq C,$
it follows from (2.12) and (2.15) that (2.14) holds.
Next, we consider the initial-boundary value problem for the coupled system (1.2) with mixed Dirichlet-Neumann boundary conditions,
$\left\{\begin{array}[]{l}u_{t}=\Delta\phi(u)+f(\cdot,u,v)\ \ \text{in}\ (0,T]\times\Omega,\\ v_{t}=\Delta v+g(\cdot,u,v)\ \ \text{in}\ (0,T]\times\Omega,\\ u=u_{0},\ \ v=v_{0}\ \ \text{in}\ \ \{0\}\times\Omega,\\ \phi(u)=\phi(u^{D})\ \text{on}\ (0,T]\times\Gamma_{1},\ \ \partial_{\nu}\phi(u)=0\ \ \text{on}\ (0,T]\times\partial\Omega/\Gamma_{1},\\ v=v^{D}\ \ \text{on}\ (0,T]\times\Gamma_{2},\ \ \partial_{\nu}v=0\ \ \text{on}\ (0,T]\times\partial\Omega/\Gamma_{2},\end{array}\right.$ (2.16)
where $\Gamma_{1},\Gamma_{2}\subset\partial\Omega$ have strictly positive Hausdorff measures, and we suppose for simplicity that $I=[0,1)$. This is motivated by applications such as the biofilm growth model (1.3), where solutions describe nonnegative quantities such as densities or concentrations. We suppose that the reaction functions $f,g:\Omega\times[0,1)\times[0,1]\to\mathbb{R}$ are measurable, bounded and uniformly Lipschitz continuous with respect to the last two arguments, i.e. there exists $L\geq 0$ such that
@@
Moreover, to ensure that the solutions remain nonegative and bounded, we assume that
$f(\cdot,0,v)\geq 0\ ,\ \ g(\cdot,u,0)\geq 0\ \text{ and }\ g(\cdot,u,1)\leq 1\ \ \text{in}\ \Omega\quad\forall u\in[0,1),v\in[0,1].$ (2.18)
@@
The condition on $f$ in (2.18) yields the nonnegativity of the solution $u$, which validates the choice of the interval $I=[0,1)$. It is not an open interval as assumed previously. However, the structural functions can be appropriately extended to the open interval $(-1,1)$. We then apply the well-posedness theory for (2.1) in combination with a comparison principle, which yields that solutions assume their values in $[0,1)$. Moreover, the conditions on $g$ in (2.18) ensure that $v$ takes values in the interval $[0,1]$. For further details, we refer to [21].
We introduce the spaces
$V_{i}=\{w\in H^{1}(\Omega):Tr\ w=0\ a.e.\ on\ \Gamma_{i}\},\quad i=1,2.$
Furthermore, solutions of system (2.16) are analogously defined as in Definition 2.1 for problem (2.1). We recall the well-posedness result for the coupled system (2.16), given in [21, Theorem 2.9].
@@
Assume that the function $\phi:[0,1)\to[0,\infty)$ satisfies the conditions (2.2)-(2.3) and $f,g$ satisfy (2.17)-(2.18). Moreover, the boundary data $u^{D}:\Omega\to[0,1)$ satisfies (2.8), $v^{D}:\Omega\to[0,1]$ satisfies $v^{D}\in H^{1}(\Omega)$, and the initial data $u_{0}:\Omega\to[0,1)$ satisfies (2.9) and $v_{0}:\Omega\to[0,1]$ is measurable. Then, there exists a unique solution $(u,v):\Omega_{T}\to[0,1)\times[0,1]$ of problem (2.16) in $L^{\infty}(0,T;L^{1}(\Omega))\times C([0,T];L^{1}(\Omega))$. The solution satisfies an $L^{1}$-contraction estimate and an energy inequality analogous to (2.12) and (2.13), respectively. Furthermore, if $\phi$ satisfies (2.6), then $(u,v)\in C([0,T];L^{1}(\Omega)\times L^{1}(\Omega))$.
In addition, if $u_{0}\leq 1-\theta$ for some $\theta\in(0,1)$, then $u\leq 1-\mu$ for some $\mu\in(0,1)$ depending on $\Omega,L$ and $\phi(1-\theta)$. In this case, if $\phi(u_{0})\in\phi(u^{D})+V$ and $\phi$ satisfies (2.6), then $\phi(u)\in H^{1}(0,T;L^{2}(\Omega))$.

### v2

@@ Abstract
We investigate the long-time behaviour of solutions of a class of singular-degenerate porous medium type equations in bounded domains with homogeneous Dirichlet boundary conditions. The existence of global attractors is shown under very general assumptions. Assuming, in addition, that solutions are globally Hölder continuous and the reaction terms satisfy a suitable sign condition in the vicinity of the degeneracy, we also prove the existence of an exponential attractor, which, in turn, yields the finite fractal dimension of the global attractor. Moreover, we extend the results for scalar equations to systems where the degenerate equation is coupled to a semilinear reaction-diffusion equation. The study of such systems is motivated by models for biofilm growth.
@@ 1.1 Singular-degenerate porous medium type equations
complemented with a suitable initial data and homogeneous Dirichlet boundary conditions, where $T>0$ and $\Omega\subset\mathbb{R}^{n},n\geq 1$, is a bounded domain. The solution $u$ assumes its values in the interval (1,1)(-1,1), $f:\Omega\times(-1,1)\to\mathbb{R}$ is continuous and $\phi:(-1,1)\to\mathbb{R}$ is strictly increasing with a degeneracy $\phi^{\prime}(0)=0$ and singularities $\phi(\pm 1)=\pm\infty$. Since the degeneracy in $u=0$ is of the same type as in the porous medium equation, $u_{t}=\Delta u^{m},m>1$, we call (1.1) a singular-degenerate equation of porous medium type. Porous medium type equations are used to model various natural phenomena such as the flow of an ideal gas through a porous medium ([2]), nonlinear heat transfer ([3]) and the spreading of certain biological populations ([4]).
@@
equipped with suitable initial values and homogeneous Dirichlet boundary data. Such systems arise in models for the growth of bacterial biofilms. Biofilms are dense aggregations of microorganisms that form in moist environments. In a biofilm, cells are encased in a slimy matrix of extracellular polymeric substances and adhere to one another and usually to surfaces. A reaction-diffusion model that predicts heterogeneous spatial structures of fully grown biofilms on the mesoscale was introduced and numerically studied by Eberl et al. in [5]. It is a system of a quasilinear reaction-diffusion equation for the biomass density $M$ and a semilinear reaction-diffusion equation for the nutrient concentration $C$. The equations are coupled via Monod reaction functions modeling biomass production and nutrient consumption. For the dimensionless variables $C$ and $M$, which are normalized, respectively, with respect to the bulk concentration and the maximum biomass density, the model reads as follows:
@@
$\phi(u)=\int\limits_{0}^{u}\frac{z^{b}}{(1-z)^{a}}dz,$
we observe that the biofilm-growth model (1.3) is a particular case of the system (1.2) that we consider. Moreover, in simulation studies, $\Omega$ is typically rectangular and mixed Dirichlet-Neumann boundary conditions are imposed. More specifically, homogeneous Neumann boundary conditions for $C$ and $M$ are assumed on the bottom part of the boundary, due to the impermeability of the surface to nutrients and biomass, and also on the lateral boundaries. Nutrients are added through the top boundary, which is described by Dirichlet boundary conditions for $C$, while homogeneous Dirichlet boundary conditions are imposed for $M$. Initially, there are small pockets of biomass of density $M_{0}<1$ on the bottom part of the boundary, resembling the initial biofilm colonies on the surface, while $M_{0}$ vanishes everywhere else in the domain. The nutrient concentration $C_{0}$ is set to the bulk concentration everywhere. The well-posedness of scalar equations and coupled systems of the form (1.1) and (1.2) with mixed Dirichlet-Neumann boundary conditions on bounded Lipschitz domains has been established in [1]. In this paper, to prove the existence of global attractors we consider smooth non-degenerate approximations and therefore consider Problems (1.1) and (1.2) with homogeneous Dirichlet boundary conditions.
@@ 1.2 Global and exponential attractors: previous results
The first existence proof for exponential attractors in [15] was non-constructive, based on the squeezing property of the semigroup and restricted to a Hilbert space setting. Efendiev et al. [17] provided an alternative, explicit construction of exponential attractors for semigroups in Banach spaces based on the decomposition the semigroup into a compact and contracting part. In [18], Czaja and Efendiev extended this construction and the existence proof for exponential attractors to a setting that can be applied to degenerate parabolic equations. It covers a noteworthy smoothing estimate formulated in suitable function spaces related to pieces of trajectories of the semigroup and not in the phase space. Efendiev and Zelik further generalized this approach allowing for estimates and function spaces that depend on points in a compact, positively invariant set to construct exponential attractors for equations of porous medium type, see [19, Remark 3.2]. For an overview and comparison of commonly used construction methods for exponential attractors we refer to [20].
@@
in $(0,\infty)\times\mathbb{R}^{n}$ and proved, by applying the splitting method, that the generated semigroup possesses a global attractor in the weighted space $L^{1}(\mathbb{R}^{n};\rho(x)dx)$ where the weight $\rho\in L^{1}(\mathbb{R}^{n})$ goes to zero at infinity. To prove the asymptotic compactness of the semigroup for our problem we use arguments from this paper. Andreu et al. [25] proved the existence of a global attractor in $L^{\infty}(\Omega)$ for (1.4) with $g=0$ in a bounded smooth domain $\Omega\subset\mathbb{R}^{n}$ with nonlinear Neumann boundary conditions under stronger assumptions on the nonlinearities $\phi$ and $f$. Moreover, Carvalho et al. [26] considered (1.4) with $\phi(u)=|u|^{m}$, $m>1$, and $g=0$ in bounded smooth domains with homogeneous Dirichlet boundary conditions and a globally Lipschitz continuous function $f$ and proved the existence of a global attractor in $H^{-1}(\Omega)$ based on the theory of maximal monotone operators. Finally, in the aforementioned paper [21], the authors showed that the biofilm growth model (1.3) in a smooth bounded domain $\Omega$ generates a semigroup which is Lipschitz continuous in $L^{1}(\Omega)\times L^{1}(\Omega)$ that possesses a global attractor.
@@ 1.3 Aim and outline
Our aim is to prove the existence of global and exponential attractors for problems (1.1) and (1.2) under suitable structural assumptions. The existence of exponential attractors ensures the existence and finite fractal dimension of the global attractor. However, we state a separate existence theorem for global attractors as it holds under much weaker assumptions. It generalizes the results in [21] on the existence of a global attractor for the biofilm model (1.3). To construct exponential attractors we additionally need a sign condition on $f$ at zero and the global Hölder continuity of solutions, similarly as in [19]. We simplify the latter proof, where first the finite fractal dimension of the global attractor was shown and then the exponential attractor was constructed, by using an abstract existence theorem for exponential attractors, see [19, Remark 3.2] and [18]. The finite fractal dimension of the global attractor is then a consequence of the existence of an exponential attractor. Our results are formulated for a broad class of degenerate problems in bounded domains with homogeneous Dirichlet boundary conditions. To the best of our knowledge, they are novel in the context of global and exponential attractors. We remark that system (1.2) is not restricted to biofilm models. Such systems can also be used to describe other phenomena with sharp interfaces propagating at a finite speed. Moreover, our results easily extend to, e.g., the model for quorum sensing induced detachment in biofilms [29] which was analytically studied in [30]. It is an extension of the biofilm growth model (1.3) involving one singular-degenerate equation and three semi-linear reaction diffusion equations for dissolved substrates.
We organize the paper as follows: In Section 2, we introduce the functional setting and state the main results on the existence of global and exponential attractors. In Section 3, we first prove the existence of a global attractor for the scalar Problem (1.1). To this end, we consider smooth nondegenerate approximations for Problem (1.1) that possess smooth classical solutions. We derive uniform bounds and smoothing estimates for the approximate solutions and passing to the limit we obtain a solution of the degenerate problem. This allows us to prove that the original problem (1.1) generates a Lipschitz continuous semigroup in a suitable phase space endowed with the $L^{1}(\Omega)$-topology. We show the existence of a bounded absorbing set and the asymptotic compactness of the trajectories, and then extend the result to the coupled system (1.2). In Section 4, under the key assumptions of global Hölder continuity of solutions and appropriate sign conditions on the partial derivative of the nonlinearity $f$ in the vicinity of the degeneracy, we prove the existence of exponential attractors for (1.1). As a consequence, the global attractor has finite fractal dimension. We also extend the existence result for exponential attractors for the coupled system (1.2).
@@ 2.1 Well-Posedness Results
We recall the well-posedness results in [1] and [31] for the scalar equation (1.1) and the coupled system (1.2), and introduce the functional setting for the associated semigroups. First, we consider the following initial-boundary value problem for (1.1) with homogeneous Dirichlet boundary conditions,
$\begin{cases}u_{t}=\Delta\phi(u)+f(\cdot,u)&\text{in}\ (0,T]\times\Omega,\\ u=u_{0}&\text{in}\ \{0\}\times\Omega,\\ \phi(u)=0&\text{on}\ (0,T]\times\partial\Omega,\end{cases}$ (2.1)
where $T>0$, $\Omega\subset\mathbb{R}^{n},n\geq 1,$ is a bounded domain with boundary $\partial\Omega$. Let $\Omega_{T}=(0,T]\times\Omega$ and I=(1,1)I=(-1,1).
@@
$\phi\ \text{is surjective,}$ (2.3)
$\left\{\begin{array}[]{ll}\phi\ \text{is piece-wise continuously differentiable, }\phi^{\prime}(0)=0\\ \phi\ \text{is\ convex on}\ [0,1)\ \text{and}\ \text{concave on}\ (-1,0].\end{array}\right$. (2.6)
@@
We assume that the initial data
@@
where $\Phi(z):=\int\limits_{0}^{z}\phi(s)ds$. We note that the quantity $\int\limits_{\Omega}\Phi(u)$ represents the (absolute) energy of the solution $u$.
Now, we introduce the functional setting for problem (2.1). We consider the closed subspace $V=H^{1}_{0}(\Omega)$ of $H^{1}(\Omega)$ and denote its dual by $V^{*}=H^{-1}(\Omega)$. The solutions need to satisfy $\phi(u)\in L^{2}(0,T;V)$ to ensure that $\phi(u)=0$ a.e. on $\partial\Omega$.
@@
$|z|\leq\delta\Phi(z)+\beta(\frac{1}{\delta}),\quad z\in I,\delta>0,$
@@
We denote by $(\cdot,\cdot)$ the dual pairing between $V^{*}$ and $V$ and by $\left$$\cdot$$\cdot\right$> the inner product in $L^{2}(\Omega)$, and define a weak solution to problem (2.1) as follows:
Definition 2.1.
@@
$\int\limits_{0}^{T}[(u_{t},\eta)+\langle\nabla\phi(u),\nabla\eta\rangle]dt=\int\limits_{0}^{T}\langle f(\cdot,u),\eta\rangle dt$ (2.9)
@@
Property (2.7), the essential boundedness of $f(\cdot,0)$ and the fact that $u\in(-1,1)$ imply that $f(\cdot,u)\in L^{\infty}(0,T;L^{2}(\Omega))$ so that the right-hand side of (2.9) is well-defined.
We formulate the following well-posedness result for the scalar problem (2.1), which follows from [[1], Theorems 2.6–2.7] in the special case of homogeneous Dirichlet boundary conditions.
@@
$\begin{split}&\lVert\Phi(u)\rVert_{L^{\infty}(0,T;L^{1}(\Omega))}+\lVert\nabla\phi(u)\rVert_{L^{2}(\Omega_{T})}^{2}\\ &\leq C\left(\lVert\Phi(u_{0})\rVert_{L^{1}(\Omega)}+\lVert f(\cdot,u)\rVert_{L^{2}(\Omega_{T})}\lVert\phi(u)\rVert_{L^{2}(\Omega_{T})}\right),\end{split}$ (2.10)
where $C\geq 0$ is some constant.
The solution depends continuously on initial data. More specifically, if $u$ and $\tilde{u}$ are two solutions of (2.1) corresponding to the functions $f$, $\tilde{f}$ satisfying (2.7) and initial data $u_{0}$, $\tilde{u_{0}}$ satisfying (2.8), then the following $L^{1}$-contraction holds,
$\lVert u(t)-\tilde{u}(t)\rVert_{L^{1}(\Omega)}\leq e^{Lt}\left(\lVert u_{0}-\tilde{u_{0}}\rVert_{L^{1}(\Omega)}+\int\limits_{0}^{t}\lVert f(\cdot,u)-\tilde{f}(\cdot,u)\rVert_{L^{1}(\Omega)}dt\right),$ (2.11)
@@
Moreover, if $\phi(u_{0})\in L^{\infty}(\Omega)$, then $\phi(u)$ is bounded by a constant depending on $\Omega$, $L$, $\lVert f(\cdot,0)\rVert_{L^{\infty}(\Omega)}$ and $\phi(u_{0})$. If the condition (2.6) holds, then we have $u\in C([0,T];L^{1}(\Omega))$. If both assumptions hold and $\phi(u_{0})\in V$, then $\phi(u)\in H^{1}(0,T;L^{2}(\Omega))$.
Next, we consider the initial-boundary value problem for the coupled system (1.2). As in the scalar case, we impose homogeneous Dirichlet boundary conditions,
$\begin{cases}u_{t}=\Delta\phi(u)+f(\cdot,u,v)&\text{in}\ (0,T]\times\Omega,\\ v_{t}=\Delta v+g(\cdot,u,v)&\text{in}\ (0,T]\times\Omega,\\ u=u_{0},\ \ v=v_{0}&\text{in}\ \ \{0\}\times\Omega,\\ \phi(u)=0,v=0&\text{on}\ (0,T]\times\partial\Omega,\end{cases}$ (2.12)
and assume that $I=[0,1)$. This is motivated by applications such as the biofilm growth model (1.3), where solutions describe nonnegative quantities such as densities or concentrations. We suppose that the reaction functions $f,g:\Omega\times[0,1)\times[0,1]\to\mathbb{R}$ are measurable, bounded and uniformly Lipschitz continuous with respect to the last two arguments, i.e. there exists $L\geq 0$ such that
@@
Moreover, to ensure that the solutions remain nonnegative and bounded, we assume that
[…]
@@
The condition on $f$ in (2.14) yields the nonnegativity of the solution $u$, which validates the choice of the interval $I=[0,1)$. It is not an open interval as assumed previously. However, the structural functions can be appropriately extended to the open interval (1,1)(-1,1). We then apply the well-posedness theory for (2.1) in combination with a comparison principle, which yields that solutions assume their values in $[0,1)$. Moreover, the conditions on $g$ in (2.14) ensure that $v$ takes values in the interval $[0,1]$. For further details, we refer to [1].
Solutions of system (2.12) are analogously defined as in Definition 2.1 for problem (2.1). We recall the well-posedness result for the coupled system (2.12), given in [1, Theorem 2.9], which we formulate for the specific case of homogeneous Dirichlet boundary conditions.
@@
Assume that the function $\phi:[0,1)\to[0,\infty)$ satisfies the conditions (2.2)-(2.3) and $f,g$ satisfy (2.13)-(2.14). Moreover, the initial data $u_{0}:\Omega\to[0,1)$ satisfies (2.8) and $v_{0}:\Omega\to[0,1]$ is measurable. Then, there exists a unique solution $(u,v):\Omega_{T}\to[0,1)\times[0,1]$ of problem (2.12) in $L^{\infty}(0,T;L^{1}(\Omega))\times C([0,T];L^{1}(\Omega))$. The solution satisfies an $L^{1}$-contraction estimate and an energy inequality analogous to (2.10) and (2.11), respectively. Furthermore, if $\phi$ satisfies (2.6), then $(u,v)\in C([0,T];L^{1}(\Omega)\times L^{1}(\Omega))$.
In addition, if $u_{0}\leq 1-\theta$ for some $\theta\in(0,1)$, then $u\leq 1-\mu$ for some $\mu\in(0,1)$ depending on $\Omega,L$ and $\phi(1-\theta)$. In this case, if $\phi(u_{0})\in V$ and $\phi$ satisfies (2.6), then $\phi(u)\in H^{1}(0,T;L^{2}(\Omega))$.